$1200 is compounded daily for 3.5 years at a rate of 8.5%. How much is in the account at the end of the term? (assume 360 days in a year)
thank you!

Answer by rfer(16322)

A=P(1+r/360)^360*3.5
A=1200(1.000236111)^1260
A=1200(1.34644)
A=$1615.73